# Adaptive Off-Policy Inference for M-Estimators Under Model Misspecification

### v1

@@ B.1 Proof of Theorem 1
Fix any $\epsilon>0$. Note that in general $\mathbbm{1}_{|Z_{t}|\geq\epsilon}>\frac{Z_{t}^{2}}{\epsilon^{2}}$. We therefore have that,
$\frac{1}{T}\sum_{t=1}^{T}\mathbb{E}\left[c^{T}\hat{V}_{t}^{-1/2}s_{t,\theta^{\star}}s_{t,\theta^{\star}}^{T}\hat{V}_{t}^{-1/2}c\mathbbm{1}_{|c^{T}\frac{1}{\sqrt{T}}\hat{V}_{t}^{-1/2}s_{t,\theta^{\star}}|\geq\epsilon}\right]$ $=\frac{1}{\epsilon^{2}T^{2}}\sum_{t=1}^{T}\mathbb{E}\left[\left(c^{T}\hat{V}_{t}^{-1/2}s_{t,\theta^{\star}}s_{t,\theta^{\star}}^{T}\hat{V}_{t}^{-1/2}c\right)^{2}\right]$
$=\frac{1}{\epsilon^{2}T^{2}}\left(\left\lVert c\right\rVert^{2}+c^{T}\left(\frac{1}{T}\sum_{t=1}^{T}A_{t}\right)c\right).$
This quantity is $o_{p}(1)$ by the same argument given above. ∎
@@
We note that by 3, there exists some $\delta_{2}>0$ such that whenever $\left\lVert\hat{\theta}_{T}-\theta^{\star}\right\rVert\geq\epsilon$, $\left\lVert\mathbb{E}\left[m_{\hat{\theta}_{T}}(X_{t},A_{t},Y_{t}\right]\right\rVert_{1}>\delta_{2}$. Therefore,
@@
However, we know that for all $\theta\in\Theta$ that
@@
By assumption, $\frac{1}{T}\sum_{t=1}^{T}\hat{V}_{t}^{-1/2}s_{t,\hat{\theta}_{T}}=o_{p}(1)$. Therefore,
@@
$\sup_{\theta\in\Theta}\left\lVert\frac{1}{T}\left(\sum_{t=1}^{T}\hat{V}_{t}^{-1/2}s_{t,\theta}-\mathbb{E}\left[\hat{V}_{t}^{-1/2}s_{t,\theta}|\mathcal{H}_{t-1}\right]\right)\right\rVert_{1}\rightarrow 0.$
We show this is true for each component individually. Define $g_{\theta}(x,a,y):=e_{j}^{T}\hat{V}_{t}^{-1/2}\dot{m}_{\theta}(x,a,y)$. By 5, for any $\epsilon>0$, $N_{[]}(\epsilon,\dot{\mathcal{M}}_{\Theta},L_{2}(\mathcal{P},\pi_{e}))<\infty$. However, because $\hat{V}_{t}^{-1/2}$ has bounded eigenvalues by assumption and $e_{j}$ is a unit vector, $N_{[]}(\epsilon,\{g_{\theta}(x,a,y):\theta\in\Theta\},L_{2}(\mathcal{P},\pi_{e}))<\infty$. Now note that
$\hat{V}_{t}^{-1/2}s_{t},\theta=\sum_{a=1}^{K}\pi_{e}(A_{t}=a|X_{t})\left(g_{\theta}(a,X_{t},f_{t}(a,X_{t}))+\mathbbm{1}_{A_{t}=a}\frac{g_{\theta}(X_{t},A_{t},Y_{t})-g_{\theta}(X_{t},A_{t},f_{t}(X_{t},A_{t}))}{\mathbb{P}\left(A_{t}=a|X_{t},\mathcal{H}_{t-1}\right)}\right),$
@@ B.2 Proof of Lemma 1
$=\underbrace{\frac{\pi_{e}(A_{t}|X_{t})}{\mathbb{P}\left(A_{t}|X_{t},\mathcal{H}_{t-1}\right)}g_{\theta}(X_{t},A_{t},Y_{t})}_{{}_{R_{t}^{(1)}(\theta)}}+\underbrace{\pi_{e}(A_{t}=a|X_{t})\sum_{a\in\mathcal{A}}\left(1-\frac{\mathbbm{1}_{A_{t}=a}}{\mathbb{P}\left(A_{t}=a|X_{t},\mathcal{H}_{t-1}\right)}\right)g_{\theta}(X_{t},a,f_{t}(X_{t},a))}_{R_{t}^{(2)}(\theta)}.$
@@
We have by 5 that for any $\delta>0$, there exists a set of brackets $B_{\delta}$ of finite size such that:
@@
    There exists a pair of functions $(l,u)\in B_{\delta}$ such that $l(x,a,y)\leq g_{\theta}(x,a,y)\leq u(x,a,y)$ for every $a\in\mathcal{A}$, $(x,y)$ in the support of $\mathcal{P}$;
@@
    $\mathbb{E}_{p_{e}}\left[u(X_{t},A_{t},Y_{t})-l(X_{t},A_{t},Y_{t})\right]\leq\delta$;
@@
$\leq\max_{(l,u)\in B_{\delta}}\frac{\pi_{e}(A_{t}|X_{t})}{\mathbb{P}\left(A_{t}|X_{t},\mathcal{H}_{t-1}\right)}u(X_{t},A_{t},Y_{t})-\mathbb{E}\left[\frac{\pi_{e}(A_{t}|X_{t})}{\mathbb{P}\left(A_{t}|X_{t},\mathcal{H}_{t-1}\right)}l(X_{t},A_{t},Y_{t})|\mathcal{H}_{t-1}\right]$
$=\max_{(l,u)\in B_{\delta}}\frac{\pi_{e}(A_{t}|X_{t})}{\mathbb{P}\left(A_{t}|X_{t},\mathcal{H}_{t-1}\right)}u(X_{t},A_{t},Y_{t})-\mathbb{E}\left[\frac{\pi_{e}(A_{t}|X_{t})}{\mathbb{P}\left(A_{t}|X_{t},\mathcal{H}_{t-1}\right)}l(X_{t},A_{t},Y_{t})|\mathcal{H}_{t-1}\right]$
$+\mathbb{E}\left[\frac{\pi_{e}(A_{t}|X_{t})}{\mathbb{P}\left(A_{t}|X_{t},\mathcal{H}_{t-1}\right)}u(X_{t},A_{t},Y_{t})|\mathcal{H}_{t-1}\right]-\mathbb{E}\left[\frac{\pi_{e}(A_{t}|X_{t})}{\mathbb{P}\left(A_{t}|X_{t},\mathcal{H}_{t-1}\right)}u(X_{t},A_{t},Y_{t})|\mathcal{H}_{t-1}\right]$
$=\max_{(l,u)\in B_{\delta}}\frac{\pi_{e}(A_{t}|X_{t})}{\mathbb{P}\left(A_{t}|X_{t},\mathcal{H}_{t-1}\right)}u(X_{t},A_{t},Y_{t})-\mathbb{E}\left[\frac{\pi_{e}(A_{t}|X_{t})}{\mathbb{P}\left(A_{t}|X_{t},\mathcal{H}_{t-1}\right)}u(X_{t},A_{t},Y_{t})\right]$
$+\mathbb{E}_{\mathcal{P},\pi_{e}}\left[u(X_{t},A_{t},Y_{t})-l(X_{t},A_{t},Y_{t})\right]$
@@
$\max_{(l,u)\in B_{\delta}}w_{t}u(X_{t},A_{t},Y_{t})-\mathbb{E}\left[w_{t}u(X_{t},A_{t},Y_{t})|\mathcal{H}_{t-1}\right]\leq\sum_{(l,u)\in B_{\delta}}|w_{t}u(X_{t},A_{t},Y_{t})-\mathbb{E}\left[w_{t}u(X_{t},A_{t},Y_{t})|\mathcal{H}_{t-1}\right]|$
@@
$\frac{1}{T}\sum_{t=1}^{T}R^{(1)}_{t}(\theta)-\mathbb{E}[R^{(1)}_{t}(\theta)|\mathcal{H}_{t-1}]\leq\delta+\frac{1}{T}\sum_{t=1}^{T}\sum_{(l,u)\in B_{\delta}}|w_{t}u(X_{t},A_{t},Y_{t})-\mathbb{E}\left[w_{t}u(X_{t},A_{t},Y_{t})|\mathcal{H}_{t-1}\right]|$
Since $|B_{\delta}|$ is finite, it is sufficient to show that $w_{t}u(X_{t},A_{t},Y_{t})-\mathbb{E}\left[w_{t}u(X_{t},A_{t},Y_{t})\right]=o_{p}(1)$ for all $(l,u)\in B_{\delta}$. To do this, we invoke the martingale weak law of large numbers 1. The first two conditions are trivially satisfied, so it is sufficient to show that $E_{\mathcal{P},p_{t}}[w_{t}^{2}u(X_{t},A_{t},Y_{t})^{2}]$ is finite for all $t$ to demonstrate the third criteria. For this, note that because of 4, $\frac{\pi_{e}(A_{t}|X_{t})}{\mathbb{P}\left(A_{t}|X_{t},\mathcal{H}_{t-1}\right)}\leq\frac{1}{\mathbb{P}\left(A_{t}|X_{t},\mathcal{H}_{t-1}\right)}<C_{1}$ so that:
$\mathbb{E}[w_{t}^{2}u(X_{t},A_{t},Y_{t})^{2}]$ $=\mathbb{E}[w_{t}^{2}u(X_{t},A_{t},Y_{t})^{2}]$
$\leq\mathbb{E}[C_{1}w_{t}u(X_{t},A_{t},Y_{t})^{2}]$
@@
$=C_{1}\mathbb{E}[\mathbb{E}_{\mathcal{P},p_{e}}[u(X_{t},A_{t},Y_{t})^{2}]]$
$=C_{1}\mathbb{E}_{\mathcal{P},p_{e}}[u(X_{t},A_{t},Y_{t})^{2}]<\infty.$
Therefore, the martingale LLN applies and we can conclude that $\frac{1}{T}\sum_{t=1}^{T}R^{(1)}_{t}(\theta)-\mathbb{E}[R^{(1)}_{t}(\theta)|\mathcal{H}_{t-1}]\leq\delta+o_{p}(1)$. Taking $\delta\rightarrow 0$ concludes the proof.
@@
We have that $p_{e}{}(A_{t}|X_{t})<1$ and $|\mathcal{A}|<\infty$ so it is sufficient to show that for all $a\in\mathcal{A}$
@@
However, we have by 6 that
$\sup_{\theta\in\Theta}\frac{1}{T}\sum_{t=1}^{T}\left(1-\frac{\mathbbm{1}_{A_{t}=a}}{\mathbb{P}\left(A_{t}=a|X_{t},\mathcal{H}_{t-1}\right)}\right)g_{\theta}(X_{t},a,f_{t}(X_{t},a))\leq\frac{1}{T}\sum_{t=1}^{T}\left(1-\frac{\mathbbm{1}_{A_{t}=a}}{\mathbb{P}\left(A_{t}=a|X_{t},\mathcal{H}_{t-1}\right)}\right)u_{\mathcal{F}}(X_{t},a,f_{t}(X_{t},a))$
Moreover,
$\mathbb{E}\left[\left(1-\frac{\mathbbm{1}_{A_{t}=a}}{\mathbb{P}\left(A_{t}=a|X_{t},\mathcal{H}_{t-1}\right)}\right)u_{\mathcal{F}}(X_{t},a,f_{t}(X_{t},a))\right]=0$
Therefore, it is again sufficient to show that $\mathbb{E}\left[\left(1-\frac{\mathbbm{1}_{A_{t}=a}}{\mathbb{P}\left(A_{t}=a|X_{t},\mathcal{H}_{t-1}\right)}\right)^{2}u_{\mathcal{F}}(X_{t},a,f_{t}(X_{t},a))^{2}\right]$ is finite for all $t$ in order for 1 to be applied. However, we note that by 4 and the fact that $\pi_{e}(A_{t}=a|X_{t})<1$ that $\left(1-\frac{\mathbbm{1}_{A_{t}=a}}{\mathbb{P}\left(A_{t}=a|X_{t},\mathcal{H}_{t-1}\right)}\right)^{2}<(1-C_{2})^{2}$. Therefore,
which is finite by assumption.

### v2

@@ B.1 Proof of Theorem 1
Fix any $\epsilon>0$. Note that in general $\mathbbm{1}_{|Z_{t}|\geq\epsilon}\leq\frac{Z_{t}^{2}}{\epsilon^{2}}$. We therefore have that,
$\frac{1}{T}\sum_{t=1}^{T}\mathbb{E}\left[c^{T}\hat{V}_{t}^{-1/2}s_{t,\theta^{\star}}s_{t,\theta^{\star}}^{T}\hat{V}_{t}^{-1/2}c\mathbbm{1}_{|c^{T}\frac{1}{\sqrt{T}}\hat{V}_{t}^{-1/2}s_{t,\theta^{\star}}|\geq\epsilon}|\mathcal{H}_{t-1}\right]$ $\leq\frac{1}{\epsilon^{2}T^{2}}\sum_{t=1}^{T}\mathbb{E}\left[\left(c^{T}\hat{V}_{t}^{-1/2}s_{t,\theta^{\star}}s_{t,\theta^{\star}}^{T}\hat{V}_{t}^{-1/2}c\right)^{2}|\mathcal{H}_{t-1}\right]$
$=\frac{1}{\epsilon^{2}T^{2}}\sum_{t=1}^{T}c^{T}\hat{V}_{t}^{-1/2}\mathbb{E}\left[(c^{T}s_{t,\theta^{\star}})^{2}(c^{T}\hat{V}_{t}^{-1/2}s_{t,\theta^{\star}})^{2}|\mathcal{H}_{t-1}\right].$
It is sufficient to show that $\mathbb{E}\left[(c^{T}s_{t,\theta^{\star}})^{2}(c^{T}\hat{V}_{t}^{-1/2}s_{t,\theta^{\star}})^{2}|\mathcal{H}_{t-1}\right]$ is bounded. By Cauchy-Schwarz, we have:
$\mathbb{E}\left[(c^{T}s_{t,\theta^{\star}})^{2}(c^{T}\hat{V}_{t}^{-1/2}s_{t,\theta^{\star}})^{2}|\mathcal{H}_{t-1}\right]$ $\leq\mathbb{E}\left[(c^{T}s_{t,\theta^{\star}})^{4}|\mathcal{H}_{t-1}\right]^{1/2}\mathbb{E}\left[(c^{T}\hat{V}_{t}^{-1/2}s_{t,\theta^{\star}})^{4}|\mathcal{H}_{t-1}\right]^{1/2}$
$\leq 2c^{4}\mathbb{E}\left[\left\lVert s_{t,\theta^{\star}}\right\rVert^{4}|\mathcal{H}_{t-1}\right]\left\lVert\hat{V}_{t}^{-1/2}\right\rVert_{\text{op}}^{2}.$
By assumption, we know that $\|\widehat{V}_{t}^{-1/2}\|_{\mathrm{op}}^{2}$ is bounded. Therefore, it is sufficient to demonstrate that $\mathbb{E}\!\left[\|s_{t,\theta^{\star}}\|^{4}\mid\mathcal{H}_{t-1}\right]=O_{p}(1)$ to conclude the proof. Recall that
$s_{t,\theta^{\star}}=\sum_{a=1}^{K}\pi_{e}(a\mid X_{t})\left\{\dot{m}_{\theta^{\star}}(X_{t},a,f_{t}(X_{t},a))+\mathbbm{1}_{\{A_{t}=a\}}\frac{\dot{m}_{\theta^{\star}}(X_{t},a,Y_{t})-\dot{m}_{\theta^{\star}}(X_{t},a,f_{t}(X_{t},a))}{\mathbb{P}(A_{t}=a\mid X_{t},\mathcal{H}_{t-1})}\right\}.$
Define
$U_{t,a}:=\dot{m}_{\theta^{\star}}(X_{t},a,f_{t}(X_{t},a)),\qquad V_{t,a}:=\mathbbm{1}_{\{A_{t}=a\}}\frac{\dot{m}_{\theta^{\star}}(X_{t},a,Y_{t})-\dot{m}_{\theta^{\star}}(X_{t},a,f_{t}(X_{t},a))}{\mathbb{P}(A_{t}=a\mid X_{t},\mathcal{H}_{t-1})}.$
Since $\pi_{e}(a\mid X_{t})\in[0,1]$ and $K<\infty$, the inequality
$\Big\|\sum_{a=1}^{K}w_{a}z_{a}\Big\|^{4}\leq K^{3}\sum_{a=1}^{K}\|z_{a}\|^{4}\quad\text{for }|w_{a}|\leq 1,$
applied with $w_{a}=\pi_{e}(a\mid X_{t})$ and $z_{a}=U_{t,a}+V_{t,a}$ yields
$\|s_{t,\theta^{\star}}\|^{4}$ $\leq K^{3}\sum_{a=1}^{K}\|U_{t,a}+V_{t,a}\|^{4}$
$\leq 8K^{3}\sum_{a=1}^{K}\big(\|U_{t,a}\|^{4}+\|V_{t,a}\|^{4}\big),$
where the second inequality follows from $(x+y)^{4}\leq 8(x^{4}+y^{4})$.
We now bound the two terms separately. For $U_{t,a}$, Assumption 6 implies that
$\|U_{t,a}\|^{4}\leq u_{F}(X_{t},a,f_{t}(X_{t},a))^{4},$
where $u_{F}$ is an envelope satisfying $\mathbb{E}[u_{F}(X_{t},a,f_{t}(X_{t},a))^{4}]<\infty$. Consequently,
$\mathbb{E}\!\left[\|U_{t,a}\|^{4}\mid\mathcal{H}_{t-1}\right]\leq\mathbb{E}\!\left[u_{F}(X_{t},a,f_{t}(X_{t},a))^{4}\mid\mathcal{H}_{t-1}\right]=O_{p}(1).$
For $V_{t,a}$, bounded importance ratios (4) imply that $\mathbb{P}(A_{t}=a\mid X_{t},\mathcal{H}_{t-1})^{-1}\leq C_{1}$, and therefore
$\|V_{t,a}\|^{4}$ $\leq C_{1}^{4}\big\|\dot{m}_{\theta^{\star}}(X_{t},a,Y_{t})-\dot{m}_{\theta^{\star}}(X_{t},a,f_{t}(X_{t},a))\big\|^{4}$
$\leq 8C_{1}^{4}\Big(\|\dot{m}_{\theta^{\star}}(X_{t},a,Y_{t})\|^{4}+\|U_{t,a}\|^{4}\Big),$
where the second line again uses $(x+y)^{4}\leq 8(x^{4}+y^{4})$. Taking conditional expectations and using the fourth-moment bound on $\dot{m}_{\theta^{\star}}(X_{t},a,Y_{t})$ from 3, together with the bound on $\mathbb{E}[\|U_{t,a}\|^{4}\mid\mathcal{H}_{t-1}]$, yields
$\mathbb{E}\!\left[\|V_{t,a}\|^{4}\mid\mathcal{H}_{t-1}\right]=O_{p}(1).$
Combining these bounds and summing over the fixed number of actions $K$ gives
$\mathbb{E}\!\left[\|s_{t,\theta^{\star}}\|^{4}\mid\mathcal{H}_{t-1}\right]\leq 8K^{3}\sum_{a=1}^{K}\left(\mathbb{E}[\|U_{t,a}\|^{4}\mid\mathcal{H}_{t-1}]+\mathbb{E}[\|V_{t,a}\|^{4}\mid\mathcal{H}_{t-1}]\right)=O_{p}(1),$
as required.
@@
We note that by 3, there exists some $\delta_{2}>0$ such that $\left\lVert\mathbb{E}\left[\dot{m}_{\hat{\theta}_{T}}(X_{t},A_{t},Y_{t})\right]\right\rVert_{1}>\delta_{2}$ will imply that $\left\lVert\hat{\theta}_{T}-\theta^{\star}\right\rVert\geq\epsilon$. Therefore,
@@
However, we know that for all $\theta\in\Theta$,
$\mathbb{E}\!\left[s_{t,\theta}\mid\mathcal{H}_{t-1}\right]$ $=\sum_{a=1}^{K}\mathbb{E}\!\Bigg[\Bigg.\pi_{e}(a\mid X_{t})\Bigg(m_{\theta}(a,X_{t},f_{t}(a,X_{t}))$
+ $\qquad\qquad+\,\mathbbm{1}_{\{A_{t}=a\}}\frac{m_{\theta}(X_{t},A_{t},Y_{t})-m_{\theta}(X_{t},A_{t},f_{t}(X_{t},A_{t}))}{\mathbb{P}(A_{t}=a\mid X_{t},\mathcal{H}_{t-1})}\Bigg)\Bigg|\mathcal{H}_{t-1}\Bigg]$
$=\mathbb{E}_{\mathcal{P},\pi_{t}}\!\left[\frac{\pi_{e}(A_{t}\mid X_{t})}{\mathbb{P}(A_{t}\mid X_{t},\mathcal{H}_{t-1})}\,m_{\theta}(X_{t},A_{t},Y_{t})\right]$
$\quad+\sum_{a=1}^{K}\mathbb{E}_{\mathcal{P},\pi_{t}}\!\left[\left(1-\frac{\mathbbm{1}_{\{A_{t}=a\}}}{\mathbb{P}(A_{t}=a\mid X_{t},\mathcal{H}_{t-1})}\right)m_{\theta}(a,X_{t},f_{t}(a,X_{t}))\right]$
$=\mathbb{E}_{\pi_{e}}\!\left[m_{\theta}(X_{t},A_{t},Y_{t})\right].$
Note that,
@@
By assumption, $\frac{1}{T}\sum_{t=1}^{T}\hat{V}_{t}^{-1/2}s_{t,\hat{\theta}_{T}}=o_{p}(1/\sqrt{T})$. Therefore,
@@
$\sup_{\theta\in\Theta}\left\lVert\frac{1}{T}\left(\sum_{t=1}^{T}\hat{V}_{t}^{-1/2}s_{t,\theta}-\mathbb{E}\left[\hat{V}_{t}^{-1/2}s_{t,\theta}|\mathcal{H}_{t-1}\right]\right)\right\rVert_{1}\overset{p}{\rightarrow}0.$
We show this is true for each component individually. Define $g_{\theta}(x,a,y):=e_{j}^{T}\hat{V}_{t}^{-1/2}\dot{m}_{\theta}(x,a,y)$. By 5, for any $\epsilon>0$, $N_{[]}(\epsilon,\dot{\mathcal{M}}_{\Theta},L_{2}(\mathcal{P},\pi_{e}))<\infty$. However, note that $\hat{V}_{t}^{-1/2}$ has bounded eigenvalues by assumption and $e_{j}$ is a unit vector, $N_{[]}(\epsilon,\{g_{\theta}(x,a,y):\theta\in\Theta\},L_{2}(\mathcal{P},\pi_{e}))<\infty$. Now note that
$e_{j}^{T}\hat{V}_{t}^{-1/2}s_{t,\theta}=\sum_{a=1}^{K}\pi_{e}(A_{t}=a|X_{t})\left(g_{\theta}(a,X_{t},f_{t}(a,X_{t}))+\mathbbm{1}_{A_{t}=a}\frac{g_{\theta}(X_{t},A_{t},Y_{t})-g_{\theta}(X_{t},A_{t},f_{t}(X_{t},A_{t}))}{\mathbb{P}\left(A_{t}=a|X_{t},\mathcal{H}_{t-1}\right)}\right),$
@@ B.2 Proof of Lemma 1
$=\underbrace{\frac{\pi_{e}(A_{t}|X_{t})}{\mathbb{P}\left(A_{t}|X_{t},\mathcal{H}_{t-1}\right)}g_{\theta}(X_{t},A_{t},Y_{t})}_{{}_{R_{t}^{(1)}(\theta)}}+\underbrace{\sum_{a=1}^{K}\pi_{e}(A_{t}=a|X_{t})\left(1-\frac{\mathbbm{1}_{A_{t}=a}}{\mathbb{P}\left(A_{t}=a|X_{t},\mathcal{H}_{t-1}\right)}\right)g_{\theta}(X_{t},a,f_{t}(X_{t},a))}_{R_{t}^{(2)}(\theta)}.$
@@
We have by assume that for any $\delta>0$, there exists a set of brackets $B_{\delta}$ of finite size. This means that for every $\theta$, there exists a pair of functions $(l,u)\in B_{\delta}$ such that:
@@
    $l(x,a,y)\leq g_{\theta}(x,a,y)\leq u(x,a,y)$ for every $a\in\mathcal{A}$, $(x,y)$ in the support of $\mathcal{P}$;
@@
    $\mathbb{E}_{\pi_{e}}\left[u(X_{t},A_{t},Y_{t})-l(X_{t},A_{t},Y_{t})\right]\leq\delta$;
@@
$\leq\max_{(l,u)\in B_{\delta}}\bigg\{\frac{\pi_{e}(A_{t}|X_{t})}{\mathbb{P}\left(A_{t}|X_{t},\mathcal{H}_{t-1}\right)}u(X_{t},A_{t},Y_{t})-\mathbb{E}\left[\frac{\pi_{e}(A_{t}|X_{t})}{\mathbb{P}\left(A_{t}|X_{t},\mathcal{H}_{t-1}\right)}l(X_{t},A_{t},Y_{t})|\mathcal{H}_{t-1}\right]\bigg\}$
$=\max_{(l,u)\in B_{\delta}}\bigg\{\frac{\pi_{e}(A_{t}|X_{t})}{\mathbb{P}\left(A_{t}|X_{t},\mathcal{H}_{t-1}\right)}u(X_{t},A_{t},Y_{t})-\mathbb{E}\left[\frac{\pi_{e}(A_{t}|X_{t})}{\mathbb{P}\left(A_{t}|X_{t},\mathcal{H}_{t-1}\right)}l(X_{t},A_{t},Y_{t})|\mathcal{H}_{t-1}\right]+$
$\qquad\mathbb{E}\left[\frac{\pi_{e}(A_{t}|X_{t})}{\mathbb{P}\left(A_{t}|X_{t},\mathcal{H}_{t-1}\right)}u(X_{t},A_{t},Y_{t})|\mathcal{H}_{t-1}\right]-\mathbb{E}\left[\frac{\pi_{e}(A_{t}|X_{t})}{\mathbb{P}\left(A_{t}|X_{t},\mathcal{H}_{t-1}\right)}u(X_{t},A_{t},Y_{t})|\mathcal{H}_{t-1}\right]\bigg\}$
$=\max_{(l,u)\in B_{\delta}}\bigg\{\frac{\pi_{e}(A_{t}|X_{t})}{\mathbb{P}\left(A_{t}|X_{t},\mathcal{H}_{t-1}\right)}u(X_{t},A_{t},Y_{t})-\mathbb{E}\left[\frac{\pi_{e}(A_{t}|X_{t})}{\mathbb{P}\left(A_{t}|X_{t},\mathcal{H}_{t-1}\right)}u(X_{t},A_{t},Y_{t})\right]+$
$\qquad\mathbb{E}_{\mathcal{P},\pi_{e}}\left[u(X_{t},A_{t},Y_{t})-l(X_{t},A_{t},Y_{t})\right]|\mathcal{H}_{t-1}\bigg\}.$
@@
$\max_{(l,u)\in B_{\delta}}w_{t}u(X_{t},A_{t},Y_{t})-\mathbb{E}\left[w_{t}u(X_{t},A_{t},Y_{t})|\mathcal{H}_{t-1}\right]\leq\sum_{(l,u)\in B_{\delta}}|w_{t}u(X_{t},A_{t},Y_{t})-\mathbb{E}\left[w_{t}u(X_{t},A_{t},Y_{t})|\mathcal{H}_{t-1}\right].$
@@
$\frac{1}{T}\sum_{t=1}^{T}R^{(1)}_{t}(\theta)-\mathbb{E}[R^{(1)}_{t}(\theta)|\mathcal{H}_{t-1}]\leq\delta+\frac{1}{T}\sum_{t=1}^{T}\sum_{(l,u)\in B_{\delta}}|w_{t}u(X_{t},A_{t},Y_{t})-\mathbb{E}\left[w_{t}u(X_{t},A_{t},Y_{t})|\mathcal{H}_{t-1}\right].$
Since $|B_{\delta}|$ is finite, it is sufficient to show that $w_{t}u(X_{t},A_{t},Y_{t})-\mathbb{E}\left[w_{t}u(X_{t},A_{t},Y_{t})|\mathcal{H}_{t-1}\right]=o_{p}(1)$ for all $(l,u)\in B_{\delta}$. To do this, we invoke the martingale weak law of large numbers from 1. Following Remark 3,it is sufficient to show that $E_{\mathcal{P},\pi_{t}}[w_{t}^{2}u(X_{t},A_{t},Y_{t})^{2}]$ is bounded by a constant for all $t$ to demonstrate the third criterion. For this, note that because of 4, there exists a constant $C_{1}>0$ such that $\frac{\pi_{e}(A_{t}|X_{t})}{\mathbb{P}\left(A_{t}|X_{t},\mathcal{H}_{t-1}\right)}\leq\frac{1}{\mathbb{P}\left(A_{t}|X_{t},\mathcal{H}_{t-1}\right)}<C_{1}$. Then
$\mathbb{E}[w_{t}^{2}u(X_{t},A_{t},Y_{t})^{2}]$ $\leq\mathbb{E}[C_{1}w_{t}u(X_{t},A_{t},Y_{t})^{2}]$
@@
$=C_{1}\mathbb{E}[\mathbb{E}_{\mathcal{P},\pi_{e}}[u(X_{t},A_{t},Y_{t})^{2}]]$
$=C_{1}\mathbb{E}_{\mathcal{P},\pi_{e}}[u(X_{t},A_{t},Y_{t})^{2}]<\infty.$
Therefore, 1 applies, and we can conclude that $\frac{1}{T}\sum_{t=1}^{T}R^{(1)}_{t}(\theta)-\mathbb{E}[R^{(1)}_{t}(\theta)|\mathcal{H}_{t-1}]\leq\delta+o_{p}(1)$. Taking $\delta\rightarrow 0$ concludes the proof.
@@
We have that $\pi_{e}{}(A_{t}|X_{t})<1$ and $|\mathcal{A}|<\infty$ so it is sufficient to show that for all $a\in\mathcal{A}$
@@
We will first show pointwise convergence and then use a covering argument to get the uniform result. Define $Z_{t,\theta}=\left(1-\frac{\mathbbm{1}_{A_{t}=a}}{\mathbb{P}\left(A_{t}=a|X_{t},\mathcal{H}_{t-1}\right)}\right)g_{\theta}(X_{t},a,f_{t}(X_{t},a))$. Note that for all $\theta$,
$\mathbb{E}[Z_{t,\theta}|\mathcal{H}_{t-1}]$ $=\mathbb{E}\left[\left(1-\frac{\mathbbm{1}_{A_{t}=a}}{\mathbb{P}\left(A_{t}=a|X_{t},\mathcal{H}_{t-1}\right)}\right)g_{\theta}(X_{t},a,f_{t}(X_{t},a))|\mathcal{H}_{t-1}\right]=0.$
By 6, $\sup_{\theta}\left\lVert g_{\theta}(X_{t},a,f_{t}(X_{t}))\right\rVert_{1}\leq u_{\mathcal{F}}(X_{t},a,f_{t}(X_{t},a))$ for all $X_{t}$ and $a$. Note that it is sufficient to show that $\mathbb{E}\left[\left(1-\frac{\mathbbm{1}_{A_{t}=a}}{\mathbb{P}\left(A_{t}=a|X_{t},\mathcal{H}_{t-1}\right)}\right)^{2}u_{\mathcal{F}}(X_{t},a,f_{t}(X_{t},a))^{2}\right]$ is bounded by a constant for all $t$ in order for 1 to be applied. However, we note that by 4 and the fact that $\pi_{e}(A_{t}=a|X_{t})<1$ that $\left(1-\frac{\mathbbm{1}_{A_{t}=a}}{\mathbb{P}\left(A_{t}=a|X_{t},\mathcal{H}_{t-1}\right)}\right)^{2}<\max((1-C_{1})^{2},1)$. Therefore,
$\mathbb{E}\left[\left(1-\frac{\mathbbm{1}_{A_{t}=a}}{\mathbb{P}\left(A_{t}=a|X_{t},\mathcal{H}_{t-1}\right)}\right)^{2}u_{\mathcal{F}}(X_{t},a,f_{t}(X_{t},a))^{2}\right]\leq\max((1-C_{1})^{2},1)\mathbb{E}\left[u_{\mathcal{F}}(X_{t},a,f_{t}(X_{t},a))^{2}\right],$
which is finite by assumption. This implies that for each $\theta$, $\frac{1}{T}\sum_{t=1}^{T}Z_{t,\theta}=0$.
Because $\Theta$ is a bounded parameter space, for any $\epsilon>0$, we can cover $\Theta\subseteq\bigcup_{j=1}^{M}\Theta_{j}$. Here, each $\Theta_{j}$ is an $\epsilon$-ball with centers denoted $\theta_{1},...,\theta_{k}$. For any $\theta$, we can decompose $Z_{t,\theta}=Z_{t,\theta}+Z_{t,\theta_{j}}$ for some $\theta_{j}$ such that $|\theta-\theta_{j}|<\epsilon$. Now, let us write out
$\sup_{\theta\in\Theta}\frac{1}{T}\sum_{t=1}^{T}Z_{t,\theta}\leq\max_{1\leq j\leq M}\lvert\frac{1}{T}\sum_{t=1}^{T}Z_{t,\theta_{j}}\rvert+\lvert\frac{1}{T}\sum_{t=1}^{T}Z_{t,\theta}-Z_{t,\theta_{j}}\rvert$.
For the first term, we know that $\max_{1\leq j\leq M}\lvert\frac{1}{T}\sum_{t=1}^{T}Z_{t,\theta}\rvert=o_{p}(1)$ because $M$ is finite for each $\epsilon$. We therefore just need to show that $\sup_{\theta\in\Theta_{j}}\lvert\frac{1}{T}\sum_{t=1}^{T}Z_{t,\theta}\rvert=o_{p}(1)$ to finish the proof.
First, note that $Z_{t,\theta}-Z_{t,\theta_{j}}$ is a martingale difference sequence. Therefore, it is sufficient to show that $\mathbb{E}\left[\left\lVert Z_{t,\theta}-Z_{t,\theta_{j}}\right\rVert^{2}\right]$ is bounded to conclude the proof. This is true becaue
$\mathbb{E}\left[\left\lVert Z_{t,\theta}-Z_{t,\theta_{j}}\right\rVert^{2}\right]$ $\leq\mathbb{E}\left[\left(1-\frac{\mathbbm{1}_{A_{t}=a}}{\mathbb{P}\left(A_{t}=a|X_{t},\mathcal{H}_{t-1}\right)}\right)^{2}\left(g_{\theta}(X_{t},a,f_{t}(X_{t},a))-g_{\theta_{j}}(X_{t},a,f_{t}(X_{t},a))\right)^{2}\right]$
$\leq\max((1-C_{1})^{2},1)\epsilon^{2}.$